# Can Graph Learning Improve Task Planning?

### v1

@@ E.2 Reformatting Details of RestBench
Type Statistic HuggingFace Multimedia Daily Life TMDB
Task Graph # Nodes 23 40 40 46
# Links 225 449 1560 979
Link Type Resource Resource Temporal Resource / Category
All Data # Samples 7,546 5,584 4,320 100
Test Set # Samples 500 500 500 94
# Avg. Nodes 3.81 3.92 4.05 2.33
# Avg. Links 2.81 2.92 3.05 1.33
Figure 8: Illustrative examples of four experimental datasets.
Appendix F Supplementary Experiments
F.1 Full Results of Training-based GNNs
In Table 8, we present comprehensive results for all training-based GNNs, including GCN [18], GAT [39], GraphSAGE [11], GIN [49], and TransformerConv [32]. To highlight the improvements brought by GNNs, we also include results from the strongest baseline, BeamSearch.
From the results, it is obvious that all GNN encoders significantly enhance the task planning abilities. For instance, when applied to Vicuna-13B on HuggingFace dataset, the introduction of GCN results in a performance improvement of $17.83\%$, GAT contributes to a $17.48\%$ increase, GraphSAGE leads to a $22.59\%$ boost, and TransformerConv improves predictions by $18.0\%$. This conclusion can be generalized across various LLMs and datasets, demonstrating the robust capabilities of GNNs.
F.2 Performance of LM+GNN Co-trained Mode
In this subsection, we conduct a supplementary study where the parameters of a pre-trained LM are also tuned along with GNN during model training. The model configuration is illustrated in Figure 7, and the results are presented in Table 9.
For LM+GNN co-trained mode, as the parameters of the LM backbone, e5-355M, are also trained, more computational resources are required. We list the number of trainable parameters and the required training times for this architecture in Table 7. In application scenarios, the choice between GNN-only and LM+GNN co-trained modes can be decided based on the availability of resources and performance requirements.
F.3 Full Efficiency Study
@@
In this subsection, we present a comprehensive efficiency study on various experimental setups: inference times across different training-free modes, time required for training GNN or LM+GNN, and the resources needed for fine-tuning LLMs. Results are shown in Tables 6 and 7.
Table 6: Efficiency Study of Training-free Methods. Due to space constraints in the table, some LLMs are abbreviated such as “GPT-3.5” for “GPT-3.5-turbo”. For Vicuna-13B and Mistral-7B, some discrepancies in reported times may be attributed to variations in deployment across different A100 devices. SGC maintains efficiency comparable to LLM’s direct inference.
@@
Mode Baichuan Vicuna CodeLlama-7B Mistral CodeLlama-13B GPT-3.5 GPT-4
@@
Table 7: Efficiency Study of Training-based Methods. We present total training times for both GNN and LM+GNN co-trained modes, and resources needed for fine-tuning LLMs.
Training GNNs
Time (Seconds)
Mode Configuration # Param HuggingFace Multimedia Daily Life
GNN-only GCN 1,049,600 136.5 136.1 237.7
GAT 1,051,648 136.9 151.8 237.8
GraphSAGE 2,098,176 134.2 149.8 233.8
GIN 2,099,200 134.2 134.9 233.6
TransformerConv 4,198,400 135.0 150.2 233.7
LM+GNN Co-trained LM+SGC 335,141,889 743.1 323.3 384.7
LM+GCN 336,191,488 482.8 567.8 384.2
LM+GAT 336,193,536 741.3 812.3 384.4
LM+GraphSAGE 337,240,064 741.2 406.8 381.7
LM+GIN 337,241,088 735.6 361.8 382.9
LM+TransformerConv 339,340,288 741.0 362.7 405.8
Fine-tuning LLMs
Device & Time (Hours)
LLM # Param (# Tunable Param) HugggingFace Multimedia Daily Life
@@
Efficiency of Training-free Modes: Open-sourced LLMs were deployed as local API services using the FastChat framework on a single A100-80G GPU. This configuration enables faster and parallel inference. Under this setup, LLM’s direct inference requires 3-15 seconds per request. GPT-3.5-turbo and GPT-4-turbo are accessed via API, with the latter generally requiring more time. GraphSearch requires several minutes to complete a request due to its exhaustive search on the task graph, impacting the efficiency. In contrast, SGC achieves comparable efficiency to LLM’s direct inference, as it requires only a single LLM query and both LM and SGC’s forward propagation processes are extremely efficient. Note that some discrepancies in reported times, such as Mistral-7B’s GreedySearch taking longer than other modes, may be attributed to variations in the deployment across different A100 GPUs.
Efficiency of Training GNNs: During experiments, each dataset shares the same GNN configuration: 1 single layer with a hidden dimension of 1024. Therefore, for each GNN, the number of parameters remains consistent across datasets. The parameter scales for GNN variants range from 1M to 4M, and the total training time for each dataset requires only 2-4 minutes, comparable to the time taken by GraphSearch to fulfill a single request. For LM+GNN co-trained mode, where e5-355M serves as the LM backbone, training times increase to approximately 6-12 minutes. In summary, both modes demonstrate high efficiency, with total training times spanning minutes, showcasing their ability to rapidly adapt to new task planning scenarios.
F.4 Case Study
Effectiveness of GNN (Qualitative Result): We show two cases in Figure 9 where the results of LLM’s direct inference, BeamSearch, and GraphSAGE are compared. Due to space constraints and issues such as LLM’s output content decoding errors or invalid paths, we present only the first four valid paths searched by BeamSearch on the task graph. From the cases, we can conclude that BeamSearch relies on LLM’s inherent reasoning abilities. Although LLM can explore the ground-truth invocation path on the task graph, their final solutions are usually not the optimal as containing either the hallucination or wrongly invoked tasks due to limited instruction following and reasoning abilities. On the contrary, GNN can effectively align decomposed steps with suitable tasks, accurately hitting the ground-truth result and enhancing task planning.
Failure Cases of GNNs: We also present the failure cases where GNN performance deteriorates compared to direct inference to provide a comprehensive discussion of our method. Our conclusion is that the method is sensitive to the quality of decomposed task steps, as an ambiguous step may mislead GNN to wrongly select the task, and such errors can cascade due to the sequential selection of tasks on the task graph. As the case shown in Figure 10, step 2 is ambiguously described as “Segment the image and identify the tabular data”, which actually incorporates two distinct steps. This ambiguty causes GNN to choose the unsuitable Tabular Classification instead of the correct Image Segmentation. Since tasks are selected sequentially on the task graph, where the next task is a neighbor of the current selection, such an error can prevent the exploration of the next appropriate task, as it may not be a neighbor of the current, incorrect selection. We also present the BeamSearch explored paths and its final solution, where it hits the ground-truth result.
Figure 10: Failure cases of GNN. Our framework relies heavily on the quality of decomposed task steps. Ambiguous steps (step 2 actually incorporates two steps) may mislead GNN to select the wrong task.
  • Successful Cases: As shown in the figure, despite inaccuracies in task decomposition (GPT-4-turbo predicts an extra step compared to the three-step ground truth), its inherent reasoning abilities and the knowledge explored along the task graph can fix these mistakes, hitting the ground-truth.
  • Failure Cases: In the failure case, although GPT-4-turbo identifies the ground-truth solution during searching process, the final decision includes an incorrect task. This occurs because the final solution selection demands complex reasoning abilities of LLM, as the context contains long textual information from different aspects: the full task graph, user request, all searched paths, and related instructions (Prompt template is shown in Table 4). The demanding context and reasoning challenge exceed GPT-4’s capabilities, leading to errors.
  • Maintaining Cases: These occur when GraphSearch merely replicates the result of GPT-4-turbo’s direct inference, indicating no added benefit from exploring the task graph. In these instances, despite navigating the graph, GPT-4 fails to self-correct inaccuracies due to inherent reasoning limitations.
We emphasize that the results of GraphSearch with GPT-4-turbo, tend to “maintaining cases”. Under such conditions, even a few failures can degrade overall performance, explaining why GraphSearch does not consistently enhance performance for GPT-4-turbo.
Figure 11: Diagnosing GraphSearch with GPT-4-turbo. We provide three types of cases: (1) Successful Cases where GPT-4 can leverage inherent reasoning abilities to select the optimal invocation path. (2) Failure Cases where GPT-4 miss in the extremely long context, containing the whole task graph, all searched paths, and instructions, selecting an unsatisfactory path. (3) Maintaining Cases where the searched result is the same as direct inference result, and those wrongly predicted tasks can still not be refined even under exhaustive search.

### v2

@@ C.2 Reformatting Details of RestBench
Type Statistic HuggingFace Multimedia Daily Life TMDB UltraTool
Task Graph # Node 23 40 40 46 260
# Links 225 449 1560 979 611
Link Type Resource Resource Temporal Resource / Category Resource
All Data # Samples 7,546 5,584 4,320 100 3,527
Test Set # Samples 500 500 500 94 500
# Avg Nodes 3.81 3.92 4.05 2.33 2.38
# Avg Links 2.81 2.92 3.05 1.33 1.38
Figure 4: Illustrative details of experimental datasets.
C.3 Reformatting Details of UltraTool
Motivation: In our main experiments using TaskBench [44] and RestBench [49], we observed that strong LLMs like GPT-4-turbo already perform well. This may be attributed to two factors: (1) the training of GPT-4-turbo likely included knowledge relevent to these benchmarks, as both were released before GPT-4 and utilize popular platforms such as HuggingFace and TMDB; (2) the task graphs are relatively small, containing no more than $50$ tasks, which falls within the capabilities of LLMs. Therefore, we aim to evacuate planning performance as well as GNN’s effectiveness in a more challenging scenario, i.e., one that requires specific knowledge beyond the LLM’s training and involves larger task graphs that possibly exceed its memory and reasoning capabilities.
Constructing a Task Graph: We utilized UltraTool [19], released in March 2024, which features complex planning scenarios across thousands of tasks, including travel, tourism, and other daily life domains. The original UltraTool contains $5,824$ samples with $2,032$ available tasks across $22$ distinct domains. However, we observed that some tasks appear in only one sample, making them quite rare. Therefore, we filtered this dataset as follows: first, we considered only tasks that appeared more than $5$ times across all samples, focusing on more common tasks that cater to daily life. Next, we retained samples that incorporated these filtered tasks, ensuring that the number of invoked tasks exceeded 2 to satisfy the multi-task planning scenario. After this filtering, we identified $260$ distinct tasks. We then constructed the task graph by treating each task as a node and adding links between tasks that were invoked sequentially in the dataset. For each task, we further verified and refined its functional descriptions to ensure semantic suitability.
Reformatting Raw Data Examples: We allocated $500$ samples for testing and $3,000$ samples for GNN training. Although the original UltraTool provided decomposed steps, we found them too coarse-grained, making it difficult to align step descriptions with suitable tasks. Therefore, we employed a similar strategy by prompting GPT-4 to infer decomposed task steps that align with each invoked task. As a result, each data sample consists of the user request, corresponding steps, and ground-truth invoked tasks, ensuring high quality for GNN training.
Appendix D Supplementary Materials for Theoretical Results
D.1 Dynamic Programming
where $\mathcal{T}(i)=\{i\}\cup\{j\;|j<i\;\text{and}\;\texttt{arr}[j]<\texttt{arr}[i]\}$ denotes the set of states that can transfer to state $i$, the aggregation function $g(x,y)$ is implemented as $\max(x,y)$, and the cost $c[i][j]$ is $1$ for those candidate states that are not equal to state $i$ as adding the element leads to a longer subsequence while 00 for the state itself.
Bellman-Ford Algorithm: The Bellman-Ford algorithm is also a classic dynamic programming algorithm used to find the shortest paths from a single source vertex to all other vertices in a weighted graph. The core idea behind the Bellman-Ford algorithm is that the distance from the source vertex to a target vertex can be computed as the minimum distance from the source to any of the target’s neighboring vertices, plus the weight of the edge connecting the neighbor to the target. Therefore, the state transition function for the Bellman-Ford algorithm can be expressed as:
where $\text{Answer}[k][i]$ represents the length of the shortest path from the source vertex to node $i$ at the $k$-th iteration, $\mathcal{T}(i)=\mathcal{N}^{-}(i)$ denotes the set of in-neighbors of node $i$, and $w[j][i]$ denotes the weight from node $j$ to $i$. The aggregation function $g(x,y)$ is implemented as $\min(x,y)$ as we try to find the shortest path.
Travelling Salesman Problem (TSP): This problem is defined as, given a set of cities and the distances between every pair of cities, finding the shortest possible route that visits every city exactly once and returns to the starting point. If we regard the set of already visited city $\mathcal{S}$ ending at $i$-th city as the current state, then states that can transfer to current state are those that ending city can reach $k$. Therefore, the state transition function for the TSP problem can be expressed as:
where $\text{Answer}[k][i][\mathcal{S}]$ represents the cost of the shortest tour that visits all the cities in the set $\mathcal{S}$ and ends at the $i$-th city, the aggregation function $g(x,y)$ is still implemented as $\min(x,y)$ since we aim to find the shortest path, and $w[j][i]$ denotes the distance from city $j$ to city $i$.
D.2 Proof of Theorem 1
Assumption 1.
Each function $f,g$ in (2) can be approximated by constant size MLP.
Assumption 2.
The aggregation function $\square$ in (2) is one of min, max, sum, mean.
Theorem 3.
(Expressiveness) Assume the input format is given in (1) and $f,g,\square$ in DP update (2) satisfy the assumptions 1 and 2. There exists a log-precision constant-depth and constant-width Transformer that simulates $1$ steps of DP update in (2). As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
Proof.
Block 1 - Initial State Broadcast: The goal of the first block is to broadcast the initial states from the initial state token to node tokens. (1) Use MLPs to recover the digits of the answer tokens and put them in the first placeholder if $e^{type}_{k}==0$; (2) Copy the first placeholder from answer token to its previous node token by using COPY in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos1}}_{k}-e^{\text{pos1}}_{j})^{2}<\delta\}$; (3) Broadcast the first placeholder with SUM in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{type1}}_{k}-e^{\text{type1}}_{j})^{2}+(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}]]$.
Block 2 - Edge Feature Operations: The goal of the second block is to copy the edge features from the edge feature token to the corresponding node tokens. (1) Use MLPs to recover the digits of the edge feature tokens and put them in the second placeholder if $e^{\text{type1}}==2$; (2) Copy the second placeholder from the edge feature token to the node token by using SUM in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos2}}_{k}-e^{\text{pos2}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}]]$.
Block 3 - Message Preparation: (1) Use MLPs to compute $g$ and place the results in the third placeholder. Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}],g(\text{Answer}[0][u_{i}],c[u_{i}][v_{i}])]$; (2) Use MLPs to clean up the first and second placeholder. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},g(\text{Answer}[0][u_{i}],c[u_{i}][v_{i}])]$
Block 4 - Message Passing: The goal of the fourth block is to compute $\square$ and $f$. (1) Use one or two attention heads (one for max, min, mean aggregations, and two for sum aggregations) to perform the aggregation operation. This is achieved by using MEAN or MAX or SUM for the first placeholder and setting $\mathcal{S}_{k}=\{j|(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}+(e^{\text{type2}}_{k}-e^{\text{type2}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\square_{v_{j}\in\mathcal{T}(u_{i})}g(\text{Answer}[0][u_{i}],c[u_{i}][v_{j}])]$; (2) Use MLPs to compute $f$. Now the state for every node token is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},f(\square_{v_{j}\in\mathcal{T}(i)}g(\text{Answer}[0][u_{i}],c[u_{i}][v_{j}]))]$.
After four blocks, the final state for every node token $u_{i}$ is given by $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[1][u_{i}]]$. $\text{Answer}[k][u_{i}]$ can be obtained by repeating the above four blocks $k$ times.
Lemma 1.
D.3 Permutation Invariance Test of LLMs
We test whether LLMs respect the permutation invariance property in graph problems and the results are given in Figure 5.
Figure 5: Illustrative Examples of LLMs Failure to Solve Graph Computational Problems under Permutation (i.e., node re-odering). Experiments were conducted for 30 times.
D.4 Proof of Proposition 1
Proposition 2.
Proof.
D.5 Proof of Theorem 2
Theorem 4.
Proof.
$-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\bm{v}_{i,j,u}\log\hat{\bm{v}}_{i,j,u}=-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\mathbb{I}_{u=v_{i,j}}\log\hat{\bm{v}}_{i,j,u}\overset{(a)}{=}-\sum_{t,v_{j-1}}\sum_{u}N_{t,v_{j-1},u}\log\hat{\bm{v}}_{i,j,u}$
where (a) uses the assumption that the output logits are determined by target node $t$ and the current node $v_{i-1}$. In (b), we assume that $\sum_{u}N_{t,v_{i-1},u}\neq 0$. The cross-entropy is minimized when $\hat{\bm{v}}_{i,j,u}=\frac{N_{t,v_{j-1},u}}{\sum_{u}N_{t,v_{j-1},u}}$. If $\sum_{u}N_{t,v_{i-1},u}=0$, then the corresponding logits will not affect the loss function and $\hat{\bm{v}}_{i,j,u}$ can take any number. ∎
Appendix E Supplementary Materials for Training-free Methods
E.1 Implementation of Baselines
In this subsection, we present the implementation details of training-free baselines, including LLM’s direct inference, GraphSearch, and ours SGC. Method illustrations are shown in Figure 6.
LLM’s Direct Inference: The prompt template for LLM’s direct inference is given in Table 4. During experiments, we uniformly apply 1-shot in context learning for LLM’s direct inference of task invocation path. For open-sourced LLMs, the temperature parameter is set to $0.2$.
GraphSearch: The prompt template for GraphSearch is given in Table 5. This algorithm conducts an iterative search on the task graph to identify an optimal task invocation path that can best satisfy a given request. In each iteration, the neighbors of the last selected task are considered as candidates. These candidates are evaluated by LLM for their suitability for the current step (Task Assessment). The search process follows a depth-first approach. After the task assessment in the final step, a set of potential invocation paths is generated. Subsequently, LLM is prompted to select the most appropriate path from these options (Path Selection). The GraphSearch algorithm is implemented in three distinct variants, each employing a unique task selection strategy:
Ours SGC: Regarding the choices of LM backbones, for integrating GPT-3.5-turbo and GPT-4-turbo with SGC, the Roberta-355M [39] serves as the text encoder. For all other datasets and LLMs, the e5-355M [61] configuration is employed.
E.2 Results of All LLMs
Table 8 supplements Table 1 with other LLMs. The proposed methods perform consistently better.
[…]
TaskBench RestBench
HuggingFace Multimedia Daily Life TMDB
SGC 56.53 29.94 2.28 56.75 31.62 2.43 62.31 36.69 3.53 32.97 9.11 1.84
SGC 59.62 31.98 2.31 61.78 37.60 2.43 83.33 63.77 3.82 48.79 15.99 1.89
SGC 63.98 39.27 2.30 67.04 45.04 2.43 87.73 70.49 3.59 66.15 42.62 1.88
SGC 67.43 42.08 2.32 74.07 49.90 2.43 87.13 66.49 3.54 64.72 25.67 1.89
SGC 65.51 39.44 2.31 73.32 53.28 2.43 92.96 79.57 3.64 71.40 47.55 1.90
SGC 76.37 50.04 2.02 83.65 63.65 2.09 96.38 86.19 3.16 82.63 59.15 1.61
SGC 77.79 52.20 2.03 88.54 69.83 2.10 97.35 85.76 3.16 82.27 56.37 1.62
E.3 Accuracy Results of Training-free Methods
[…]
LLM Method HuggingFace Multimedia DailyLife LLM Method HuggingFace Multimedia DailyLife
Vicuna 13B Direct 8.72 11.20 24.43 CodeLlama 7B Direct 15.00 15.19 47.69
GreedySearch 10.95 9.34 3.76 GreedySearch 16.20 20.04 45.07
AdaptiveSearch 10.55 10.37 13.15 AdaptiveSearch 18.79 19.41 46.48
BeamSearch 12.58 12.03 11.06 BeamSearch 17.00 21.10 45.67
SGC 16.02 20.12 42.17 SGC 21.20 29.32 55.33
Mistral 7B Direct 16.36 25.05 44.52 CodeLlama 13B Direct 14.29 24.10 66.40
GreedySearch 20.45 16.02 29.22 GreedySearch 19.11 24.90 67.00
AdaptiveSearch 21.88 26.90 49.32 AdaptiveSearch 17.30 24.10 66.80
BeamSearch 20.45 29.36 45.89 BeamSearch 19.92 25.70 67.20
SGC 25.15 33.68 52.28 SGC 22.54 36.75 70.80
GPT- 3.5-turbo Direct 28.95 47.96 81.30 GPT- 4-turbo Direct 33.68 60.56 86.77
GreedySearch 26.90 52.47 73.17 GreedySearch 33.68 61.37 86.77
AdaptiveSearch 29.36 51.61 74.59 AdaptiveSearch 33.47 61.17 86.77
BeamSearch 32.03 52.47 80.87 BeamSearch 33.26 61.57 86.57
SGC 32.44 51.61 83.13 SGC 34.09 60.97 86.77
Due to space limitations, we present only the Node-F1 and Link-F1 scores for training-free methods in the main text. Here, we provide the Accuracy results in Table 9. These results show that integrating SGC significantly enhances accuracy across different LLMs on all datasets, making previously unsolvable planning scenarios manageable and successful.
E.4 Computational Cost Analysis
@@
In this subsection, we present a comprehensive efficiency study on inference time of training-free methods and results are shown in Table 10.
Table 10: Computational Cost Analysis of Training-free Methods. Due to space constraints in the table, some LLMs are abbreviated such as “GPT-3.5” for “GPT-3.5-turbo”.
@@
Method Baichuan Vicuna CodeLlama-7B Mistral CodeLlama-13B GPT-3.5 GPT-4
@@
Open-sourced LLMs were deployed as local API services using the FastChat framework [^2] https://github.com/lm-sys/FastChat on a single A100-80G GPU. This configuration enables faster and parallel inference. Under this setup, LLM’s direct inference requires 3-15 seconds per request. GPT-3.5-turbo and GPT-4-turbo are accessed via API, with the latter generally requiring more time. GraphSearch requires several minutes to complete a request due to its exhaustive search on the task graph, impacting the efficiency. In contrast, SGC achieves comparable efficiency to LLM’s direct inference, as it requires only a single LLM query and both LM and SGC’s forward propagation processes are extremely efficient (typically completing within seconds). Note that some discrepancies in reported times, such as Mistral-7B’s GreedySearch taking longer than other modes, may be attributed to variations in the deployment across different A100 services.
Appendix F Supplementary Materials for Training-based Methods
F.1 Implementation of Training-based GNNs
Figure 7: Illustration of our GNN-enhanced Task Planning. First, LLMs interpret user request into several manageable steps. Then, we leverage GNNs for task retrieval, sequentially matching each step description to a suitable task, finally generating the invocation path.
LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-355M model [61] as the LM backbone. For the graph encoder, our setup includes a single layer with a hidden dimension of $1024$. During the model training, we set the batch size to $512$ and run for $20$ epochs with a learning rate of $31e-$ 3. We use the Adam optimizer [24] and implement an early stopping mechanism with a patience of $5$ epochs to prevent over-fitting. All experiments are conducted on a single NVIDIA A100-80G GPU.
Training Data Preparation: From each dataset in TaskBench, we randomly select $3,000$ samples to create the trainset. The original data includes specific task steps and corresponding task invocation paths. Therefore, we first employ a topological sort to align each task step accurately with corresponding task, forming “<step, ground-truth task>” pairs. Then, for each pair, we randomly sample two negative tasks to constitute the “<step, positive task, negative task>” triplets for model training. These negative samples are selected based on how textually similar they are to the positive one, creating a robust differentiation challenge for the model.
Choices of Different Configurations: Our training-based model offers two configuration options: training only the GNN while keeping the LM frozen, or co-training both the LM and GNN. Illustrations of these configurations are provided in Figure 8. The first configuration is designed to explore the GNN’s capability in task retrieval. The latter leverages the LM’s dataset-specific semantic embeddings to enhance performance. For the co-training setup, we use a learning rate of $52e-$ 5 and a training duration of $10$ epochs.
F.2 Implementation of TAPE and GraphToken
We adapt TAPE [16] and GraphToken [37] as training-required baselines for task planning. Here, we detail the adaptation processes for each method.
TAPE: To adapt TAPE for task planning, we reformulate the planning task as a node classification problem, aiming to classify a user request into the appropriate task labels within the task graph. Firstly, LLMs interpret user requests by generating high-quality explanatory text, i.e., chain-of-thought reasoning, to understand each request. Then, a fine-tuned LM encodes these chain-of-thought texts into latent embeddings. The fine-tuning process is based on pairs of textual descriptions and their corresponding ground-truth tasks. Finally, GNNs select the suitable tasks by leveraging both the generated text embeddings and the task embeddings. For a fair performance comparison, we fine-tune the LM as e5-355M and configure the GNN as a $2$-layer GraphSAGE with a hidden dimension of $1024$. The training data for both the LM and GNNs are consistent, utilizing the same split of the training dataset.
GraphToken: For adapting GraphToken, we first encode the task graph $G(V,E,X)$ by computing each node’s representation $\bm{x}_{v}$ by feeding its descriptive text into the pre-trained LM, e5-355M. Then, a GNN transforms the initial node embedding matrix into $H$. The GNN is configured as a $2$-layer GraphSAGE with a hidden dimension of $1024$. A mean pooling operation is further applied to the node embeddings to obtain the graph’s overall representation as $\bm{h}_{G}=\text{Mean-Pooling}(H)$. Finally, both $\bm{h}_{G}$ and the text embeddings of input instruction and user request are concatenated and fed into the LLM to generate the output planning result. During this process, the backbone LLMs are frozen, and only the GNN’s parameters are tuned. For each dataset, we train over $4$ epochs using the same $3,000$ training samples consisting of <user request, ground-truth planning> pairs. The batch size is set to $16$, with a maximum input length of $512$ and a maximum output length of $300$ for HuggingFace and Multimedia datasets, and $600$ for DailyLife dataset. The learning rate is configured to $51e-$ 5.
F.3 Implementation of Fine-tuning LLMs
To explore the effectiveness of our proposed framework on fine-tuned LLMs, we employ Supervised Fine-Tuning (SFT) on the CodeLlama-7B and Vicuna-13B models. We use the same set of $3,000$ samples from GNN training as the LLMs’ fine-tuning data. In fine-tuning the LLM with LoRA [18], we set the lora_r parameter (dimension for LoRA update matrices) to $8$ and the lora_alpha (scaling factor) to $16$. The dropout ratio is set to $0.1$, the batch size to $2$, and we conduct training over 2 epochs with a learning rate of $51e-$ 5. For HuggingFace dataset, the maximum input length is set to 800, while the maximum output length is 400. For Multimedia and Daily Life datasets, which contain a larger number of tasks and require longer textual inputs, we set the maximum input and output length to 1000 and 500, respectively. We utilize 2 NVIDIA A100-80G GPUs for fine-tuning the LLMs.
F.4 Full Results of Training-based GNNs
In Table 11, we present comprehensive results for all training-based GNNs, including GCN [25], GAT [56], GraphSAGE [15], GIN [66], and Graph Transformer [46]. To highlight the improvements brought by GNNs, we also include results from the strongest baseline, BeamSearch.
From the results, it is obvious that all GNN encoders significantly enhance the task planning abilities. For instance, when applied to Vicuna-13B on HuggingFace dataset, the introduction of GCN results in a performance improvement of $17.83\%$, GAT contributes to a $17.48\%$ increase, GraphSAGE leads to a $22.59\%$ boost, and GraphTransformer improves predictions by $18.0\%$. This conclusion can be generalized across various LLMs and datasets, demonstrating the robust capabilities of GNNs.